Ben has eight 5 cent coins and three 10 cent coins. How much does he have altogether?
(I need a step by step solution please)

Answer by oberobic(2304)

5 cent coins are called nickels.
10 cent coins are called dimes.
.
8 nickels = 8 * 5 cents = 40 cents
3 dimes = 3 * 10 cents = 30 cents
40 cents + 30 cents = 70 cents
.
Done.